OpenStudy (anonymous):

Worlds population in 1960 was approximately 3 billion; in 2000 it was approximately 6 billion. Model this using a differential equation of your choice. You should include your reasons for your choice of models, the solution to the equation, and a graph spanning the given data and some time into the future. Please help me get this finished!!! Medal when done! :)

OpenStudy (anonymous):

Population growth is usually modeled so that future population growth rate depends linearly on the current size, \(\frac{dP}{dt}=kP\) for some constant \(k\) where \(P\) is the population (say, in millions). a function of \(t\), say, the time in years since 1960... we're given boundary conditions, particularly \(P(0)=3\) and \(P(40)=6\.)

OpenStudy (anonymous):

Okay so when I write this equation out p(0)=3 and P(40)=3/20?

OpenStudy (tkhunny):

?? Why did the population shrink? \(\dfrac{dP}{dt} = kP\) This is separable. Solve the differential equation, then worry about the known values.

OpenStudy (anonymous):

It didn't shrink...It went from 3 billion to 6 billion.

OpenStudy (tkhunny):

Not if p(0) = 3 and p(40) = 3/20 -- Where did '6' go?